### v1

# Exact Exploratory Bi-factor Analysis: A Constraint-based Optimisation Approach

Abstract
Bi-factor analysis is a form of confirmatory factor analysis widely used in psychological and educational measurement. The use of a bi-factor model requires the specification of an explicit bi-factor structure on the relationship between the observed variables and the group factors. In practice, the bi-factor structure is sometimes unknown, in which case an exploratory form of bi-factor analysis is needed to find the bi-factor structure. Unfortunately, there are few methods for exploratory bi-factor analysis, with the exception of a rotation-based method proposed in Jennrich and Bentler, (2011, 2012). However, this method only finds approximate bi-factor structures, as it does not yield an exact bi-factor loading structure, even after applying hard thresholding. In this paper, we propose a constraint-based optimisation method that learns an exact bi-factor loading structure from data, overcoming the issue with the rotation-based method. The key to the proposed method is a mathematical characterisation of the bi-factor loading structure as a set of equality constraints, which allows us to formulate the exploratory bi-factor analysis problem as a constrained optimisation problem in a continuous domain and solve the optimisation problem with an augmented Lagrangian method. The power of the proposed method is shown via simulation studies and a real data example. Extending the proposed method to exploratory hierarchical factor analysis is also discussed. The codes are available on “https://anonymous.4open.science/r/Bifactor-ALM-C1E6”.
Keywords: Bi-factor model, augmented Lagrangian method, exploratory bi-factor analysis, hierarchical factor model
1 Introduction
The bi-factor model was originally proposed by Holzinger and Swineford, (1937) for linear factor analysis and further extended by Gibbons and Hedeker, (1992), Gibbons et al., (2007), Cai et al., (2011), among others, to nonlinear factor analysis settings to account for dichotomous, ordinal, and nominal data. These models assume the observed variables can be accounted for by $(G+1)$ factors, with a general factor, onto which all items load directly, and $G$ group factors that each is associated with a subset of variables. Such a specification leads to good interpretations in many real-world applications. These models have received wide applications in psychological and educational measurement; see (e.g., Bradlow et al.,, 1999; Cai et al.,, 2016; Chen et al.,, 2012; DeMars,, 2006, 2012; Gibbons et al.,, 2009; Gignac and Watkins,, 2013; Jeon et al.,, 2013; Reise et al.,, 2007; Rijmen,, 2010). However, we note that all these applications of bi-factor analysis are confirmatory, in the sense that one needs to pre-specify the number of group factors and the relationship between the observed variables and the group factors. Such prior knowledge may not always be available. In that case, an exploratory form of bi-factor analysis is needed.
However, standard exploratory factor analysis methods are not applicable to the bi-factor analysis setting, and few methods have been developed for exploratory bi-factor analysis. An exception is the seminal work of Jennrich and Bentler, (2011, 2012), who proposed a rotation-based method for exploratory bi-factor analysis with orthogonal and oblique factors, respectively. However, their approach has some limitations. First, as a common issue with rotation-based methods, their method does not yield many zero loadings, and thus, the resulting loading structure does not have an exact bi-factor structure. Although a post-hoc thresholding procedure (i.e., treating loadings with an absolute value below a threshold as zero) can be applied to obtain a cleaner loading pattern, it does not work well when some variables show relatively large loadings on more than one group factor after the rotation. In fact, one cannot always find a threshold that yields an exact bi-factor structure that each variable loads on one and only one group factor. Second, as noted in Jennrich and Bentler, (2012), their method fails completely in the best case where there is a rotation of an initial loading matrix that has an exact bi-factor structure. This failure is due to that their rotation method cannot incorporate the zero constraints on the correlations between the general factor and the group factors.
This paper proposes a constrained optimisation method for exploratory bi-factor analysis, which overcomes the issues with the rotation-based method. The contribution is four-fold. First, we provide a mathematical characterisation of the bi-factor loading structure as a set of nonlinear equality constraints, which allows us to formulate the exploratory bi-factor analysis problem as a constrained optimisation problem. In other words, it turns a discrete model selection problem into a continuous optimisation problem, which reduces the computational demand in some sense. It is shown that in the aforementioned best case where the rotation method fails, the global solutions to the optimisation can perfectly recover the exact bi-factor structure. Second, we propose an augmented Lagrangian method (ALM, Bertsekas,, 2014) for solving this optimisation problem, which is a standard numerical optimisation method for solving constrained optimisation with robust empirical performance and good theoretical properties. Third, we combine the proposed method with the Bayesian information criterion (BIC, Schwarz,, 1978) for selecting the number of group factors. Compared with existing exploratory factor analysis methods for determining the number of factors, our method is tailored to the bi-factor model structure and, thus, tends to be statistically more efficient when the data is indeed generated by a bi-factor model. Finally, we demonstrate that the proposed method can be extended to learning the loading structure of hierarchical factor models (Schmid and Leiman,, 1957; Yung et al.,, 1999), a higher-order extension of the bi-factor model that has received wide applications (see, e.g., Brunner et al.,, 2012, and references therein). The bi-factor model can be viewed as a special hierarchical factor model with a two-layer factor structure, with the general factor in one layer and the group factors in the other. Similar to exploratory bi-factor analysis, the proposed method yields exact hierarchical factor loading structures without a need for post-hoc treatments.
The rest of the paper is organised as follows. In Section 2, we formulate the exploratory bi-factor analysis problem as a constrained optimisation problem and propose an ALM for solving it. We also propose a BIC-based procedure for selecting the number of group factors. Simulation studies and a real data example are presented in Sections 3 and 4, respectively, to evaluate the performance of the proposed method. We conclude with discussions in Section 5. An extension of the proposed method to exploratory hierarchical factor analysis is given in the Appendix, including the extended method and a simulation study.
2 Exploratory Bi-factor Analysis by Constrained Optimisation
2.1 Bi-factor Model Structure and a Constrained Optimisation Formulation
For the ease of exposition and simplification of the notation, we focus on the linear bi-factor model, while noting that the constraints that we derive for the bi-factor loading matrix below can be combined with the likelihood function of other bi-factor models (e.g., Gibbons and Hedeker,, 1992; Gibbons et al.,, 2007; Cai et al.,, 2011) for their exploratory analysis. We focus on the extended bi-factor model, also known as the oblique bi-factor model, as considered in Jennrich and Bentler, (2012) and Fang et al., (2021). This model is more general than the standard bi-factor model, in the sense that the latter assumes all the factors to be uncorrelated while the former allows the group factors to be correlated. As established in Fang et al., (2021), this extended bi-factor model is identifiable under mild conditions. We note that the proposed method can be easily adapted to the standard bi-factor model.
Consider a dataset with $N$ observation units from a certain population and $J$ observed variables. The extended bi-factor model assumes that the $J$ variables are partitioned into $G$ non-overlapping groups. Let $\mathcal{B}_{g}\subset\{1,...,J\}$ denote the set of variables in the $g$th group. Then the bi-factor model assumption implies that $\mathcal{B}_{g}\cap\mathcal{B}_{g^{\prime}}=\emptyset$, $g\neq g^{\prime}$, and $\cup_{g=1}^{G}\mathcal{B}_{g}=\{1,...,J\}$. The model further assumes that the population covariance matrix of the observed variables can be decomposed as
$\Sigma=\Lambda\Phi\Lambda^{\top}+\Psi,$
where $\Lambda=(\lambda_{jg})_{J\times(G+1)}$ is the loading matrix, $\Phi=(\phi_{gg^{\prime}})_{(G+1)\times(G+1)}$ is the correlation matrix of the factors, which is assumed to be strictly positive definite, and $\Psi$ is a $J\times J$ diagonal matrix with diagonal entries $\psi_{1}$, …, $\psi_{J}$. The bi-factor structure imposes the following zero constraints on the loading matrix $\Lambda$:
[…]
In addition, the correlation matrix $\Phi$ satisfies $\phi_{1k}=0$ for all $k\neq 1$, meaning that all the group factors are uncorrelated with the general factor. This constraint on $\Phi$ is necessary to ensure that the extended bi-factor model is identifiable (Fang et al.,, 2021), as otherwise, there will be a rotational indeterminacy issue.
Now suppose that the number of group factors $G$ is known, while the variable partition, $\mathcal{B}_{g},g=1,...,G$, is unknown. Section 2.3 considers the selection of $G$ when it is unknown. The bi-factor structure means that for each $j$, there is at most one non-zero element in $(\lambda_{j,2},...,\lambda_{j,G+1})^{\top}$. Consequently, the loading matrix $\Lambda$ should satisfy the following $J(G-1)G/2$ constraints:
$\lambda_{jk}\lambda_{jk^{\prime}}=0,\mbox{~{}for all~{}}k,k^{\prime}=2,...,G+1,k\neq k^{\prime},j=1,...,J$ (1)
Therefore, the exploratory bi-factor analysis problem can be translated into the following constrained optimisation problem
$\min_{\Lambda,\Phi,\Psi}{}{}{}$ $l(\Lambda\Phi\Lambda^{\top}+\Psi;S)$ (2)
$s.t.{}{}{}$ $\lambda_{jk}\lambda_{jk^{\prime}}=0,\mbox{~{}for all~{}}k,k^{\prime}=2,...,G+1,k\neq k^{\prime},j=1,...,J,$
$\phi_{1k}=0,k=2,...,G+1,\Phi\mbox{~{}is correlation matrix},$
and Ψ is a diagonal $\mbox{and~{}}\Psi\mbox{~{}is a diagonal matrix},$
where $l$ is a loss function and $S$ is the sample covariance matrix of observed data. We focus on the case when $l$ is the fit function based on the normal likelihood (Chen et al.,, 2023)
$l(\Lambda\Phi\Lambda^{\top}+\Psi;S)=N\big{(}\log(\text{det}(\Lambda\Phi\Lambda^{\top}+\Psi))+\textnormal{tr}(S(\Lambda\Phi\Lambda^{\top}+\Psi)^{-1})-\log(\text{det}(S))-J\big{)},$ (3)
while noting that this loss function can be replaced by other loss functions for factor analysis (see, e.g., Chen et al.,, 2023), including the Frobeneious norm of $\Lambda\Phi\Lambda^{\top}+\Psi-S$ that is used in the least square estimator for factor analysis.
The following theorem shows that the proposed method can perfectly recover the true bi-factor loading structure in the best case when $S=\Lambda^{*}\Phi^{*}(\Lambda^{*})^{\top}+\Psi^{*}$ is the true covariance matrix of data, where $\Lambda^{*}$, $\Phi^{*}$ and $\Psi^{*}$ denote the true value of the corresponding parameter matrices. Note that the rotation method proposed in Jennrich and Bentler, (2012) completely fails in this case.
[…]
The following conditions are sufficient for the identifiability of the bi-factor structure and its parameters.
Condition 1.
Given $S=\Lambda^{*}\Phi^{*}(\Lambda^{*})^{\top}+\Psi^{*}$. Suppose that there exists another pair of parameters $\Lambda,\Phi,\Psi$ satisfy the bi-factor structure constraints, we have $\Lambda^{*}\Phi^{*}(\Lambda^{*})^{\top}=\Lambda\Phi(\Lambda)^{\top}$ and $\Psi^{*}=\Psi$.
Condition 2.
$|\mathcal{H}^{*}|\geq 2$. In addition, there exists $g_{1}\in\mathcal{H}^{*}$ such that $|Q_{g_{1}}^{*}|\geq 3$ and any 2 rows of $\Lambda^{*}[\mathcal{B}^{*}_{g_{1}},\{1,1+g_{1}\}]$ are linearly independent.
Theorem 1.
Suppose that Conditions 1-2 hold. For any parameters $\Lambda,\Phi,\Psi$ that satisfy $S=\Lambda\Phi(\Lambda)^{\top}+\Psi$, there exist a diagonal sign-flip matrix $D\in\mathcal{D}$ and a permutation matrix $P\in\mathcal{P}$ such that $\Lambda=\Lambda^{*}PD$ and $\Phi=DP^{\top}\Phi^{*}PD$.
The Proof of Theorem 1 is given in the appendix.
Remark 1.
We note that without additional information, the best we can achieve is to recover $\Lambda^{*}$ and $\Psi^{*}$ up to $\Lambda=\Lambda^{*}PD$ and $\Psi=DP^{\top}\Psi^{*}PD$, where the permutation matrix $P$ and sign-flip matrix $D$ are necessary to account for the label and sign indeterminacies of the factor model. In that case, $\Lambda^{*}\Phi^{*}(\Lambda^{*})^{\top}=\Lambda\Phi(\Lambda)^{\top}$.
Remark 2.
Condition 1 ensures the separation between low rank matrix $\Lambda^{*}\Phi^{*}(\Lambda^{*})^{\top}$ and diagonal matrix $\Psi^{*}$. One of the sufficient condition of Condition 1 that can be easily checked in practice is given in Condition 3 below. We refer to Theorem 5.1 in Anderson and Rubin, (1956) and Theorem 2 in Fang et al., (2021) for further choices of conditions to replace Condition 1.
Condition 3.
$|\mathcal{Q}^{*}_{g}|\geq 3$ for all $g=1,\ldots,G$ and $|\mathcal{H}^{*}|\geq 3$.
Remark 3.
Conditions 2 is slightly different with the Condition E3S for Proposition 1 in Fang et al., (2021), which shows the identifiability of parameters when the bi-factor structure is known. Here, combined with Condition 1, we show that such conditions are sufficient to ensure the identifiability under an unknown bi-factor structure.
2.2 Proposed ALM
Following the previous discussion, we see that we can perform exploratory bi-factor analysis by solving the optimisation problem with some equality constraints and the constraint that $\Phi$ is a correlation matrix. To deal with the constraints in $\Phi$, we consider the following reparameterisation that has been considered in Alfonzetti et al., (2024), which is also similar to the implementation in the state-of-the-art statistical software Stan (Stan Development Team,, 2022):
$\Phi=\left(\begin{bmatrix}1&\mathbf{0}^{\top}\\ \mathbf{0}&U^{T}\end{bmatrix}\right)\left(\begin{bmatrix}1&\mathbf{0}^{\top}\\ \mathbf{0}&U\end{bmatrix}\right),$ (4)
where $U$ is defined recursively by
[…]
Here $z_{ij}=\tanh(\gamma_{ij})$ is the Fisher’s transformation of $G(G-1)/2$ unconstrained parameters $\gamma_{ij}$.
With slight abuse of notation, we reexpress the covariance matrix as $\Phi(\bm{\gamma})$. In addition, we use $\bm{\psi}=(\psi_{1},...,\psi_{J})^{\top}$ to denote the vector of diagonal entries of $\Psi$ and reexpress the residual covariance matrix as $\Psi(\bm{\psi})$. Thus, the optimisation problem (2) is now simplified as
$\min_{\Lambda,\bm{\gamma},\bm{\psi}}{}{}{}$ $l(\Lambda\Phi(\bm{\gamma})\Lambda^{\top}+\Psi(\bm{\psi});S)$ (6)
$s.t.{}{}{}$ $\lambda_{jk}\lambda_{jk^{\prime}}=0,\mbox{~{}for all~{}}k,k^{\prime}=2,...,G+1,k\neq k^{\prime},j=1,...,J,$
which is an equality-constrained optimisation problem.
The standard approach for solving such a problem is the augmented Lagrangian method (e.g., Bertsekas,, 2014). This method aims to find a solution to (6) by solving a sequence of unconstrained optimisation problems. Let $t$ denote the $t$th unconstrained optimisation problem in the ALM. The corresponding objective function, also known as the augmented Lagrangian function, takes the form
$\min_{\Lambda,\bm{\gamma},\bm{\psi}}~{}~{}~{}l(\Lambda\Phi(\bm{\gamma})\Lambda^{\top}+\Psi(\bm{\psi});S)+\left(\sum_{j=1}^{J}\sum_{k=2}^{G}\sum_{k^{\prime}=k+1}^{G+1}\beta_{jkk^{\prime}}^{(t)}\lambda_{jk}\lambda_{jk^{\prime}}\right)+c^{(t)}\left(\sum_{j=1}^{J}\sum_{k=2}^{G}\sum_{k^{\prime}=k+1}^{G+1}(\lambda_{jk}\lambda_{jk^{\prime}})^{2}\right),$ (7)
where $c^{(t)}$ and $\beta_{jkk^{\prime}}^{(t)}$s are auxiliary coefficients of the ALM determined by the result of the $(t-1)$th optimisation. Details of the ALM are given in Algorithm 1 below, where the function $h$ returns the second-largest value of a vector.
Algorithm 1 Augmented Lagrangian Method for Exact Exploratory Bi-factor Analysis
1:Initial value $\Lambda^{(0)}$, $\bm{\gamma}^{(0)}$ and $\bm{\psi}^{(0)}$, initial lagrangian parameters $\beta_{j,k,k^{\prime}}^{(0)}$ for $j=1,\ldots,J$, $k=2,\ldots,G$ and $k^{\prime}=k+1,\ldots,G+1$, initial penalty coefficient $c^{(0)}>0$, constants $c_{\theta}\in(0,1)$ and $c_{\sigma}>1$, tolerances $\delta_{1},\delta_{2}>0$.
2:while $t=1,2,\ldots$ do
3:     Solve the following problem:
$\Lambda^{(t)},\bm{\gamma}^{(t)},\bm{\psi}^{(t)}$
$=$ $\arg\min_{\Lambda,\bm{\gamma},\bm{\psi}}~{}~{}~{}l(\Lambda\Phi(\bm{\gamma})\Lambda^{\top}+\Psi(\bm{\psi});S)+\left(\sum_{j=1}^{J}\sum_{k=2}^{G}\sum_{k^{\prime}=k+1}^{G+1}\beta_{jkk^{\prime}}^{(t)}\lambda_{jk}\lambda_{jk^{\prime}}\right)+c^{(t)}\left(\sum_{j=1}^{J}\sum_{k=2}^{G}\sum_{k^{\prime}=k+1}^{G+1}(\lambda_{jk}\lambda_{jk^{\prime}})^{2}\right).$
4:     if $d\left((\Lambda^{(t+1)},\bm{\gamma}^{(t+1)},\bm{\psi}^{(t+1)}),(\Lambda^{(t)},\bm{\gamma}^{(t)},\bm{\psi}^{(t)})\right)<\epsilon$ then
5:         return $\Lambda^{(t+1)},\bm{\gamma}^{(t+1)},\bm{\psi}^{(t+1)}$.
6:     end if
7:     $\beta_{jkk^{\prime}}^{(t+1)}=\beta_{jkk^{\prime}}^{(t)}+c^{(t)}\lambda^{(t)}_{jk}\lambda^{(t)}_{jk^{\prime}}$;
8:     if $\left(\sum_{j=1}^{J}\sum_{k=2}^{G}\sum_{k^{\prime}=k+1}^{G+1}(\lambda_{jk}^{(t)}\lambda_{jk^{\prime}}^{(t)})^{2}\right)^{1/2}>c_{\theta}\left(\sum_{j=1}^{J}\sum_{k=2}^{G}\sum_{k^{\prime}=k+1}^{G+1}(\lambda_{jk}^{(t-1)}\lambda_{jk^{\prime}}^{(t-1)})^{2}\right)^{1/2}$ then
9:         $c^{(t+1)}=c_{\sigma}c^{(t)}$;
10:     else
11:         $c^{(t+1)}=c^{(t)}$;
12:     end if
13:     if $\left(\|\Lambda^{(t)}-\Lambda^{(t-1)}\|_{F}^{2}+\|\bm{\gamma}^{(t)}-\bm{\gamma}^{(t-1)}\|^{2}+\|\bm{\psi}^{(t)}-\bm{\psi}^{(t-1)}\|^{2}\right)^{1/2}/\sqrt{J(G+2)+G(G-1)/2}<\delta_{1},$
14:and $\max_{j\in\{1,...,J\}}h(|\lambda_{j2}^{(t)}|,...,|\lambda_{j,G+1}^{(t)}|)<\delta_{2}$ then
15:         Break
16:     end if
17:end while
18:$\Lambda^{(t)},\bm{\gamma}^{(t)},\bm{\psi}^{(t)}$.
The updating rule of $\beta^{(t)}_{jkk^{\prime}}$ and $c^{(t)}$ follows equation (1) and (47) in Chapter 2.2 of Bertsekas, (2014) and the convergence of Algorithm 1 is guaranteed by Proposition 2.7 of Bertsekas, (2014). We follow the recommended choices of $c_{\theta}=0.25$ and $c_{\sigma}=10$ in Bertsekas, (2014).
We remark on the stopping criterion in the implementation of Algorithm 1. We monitor the convergence of the algorithm based on two criteria: (1) the change in parameter values in two consecutive steps, measured by
$\left(\|\Lambda^{(t)}-\Lambda^{(t-1)}\|_{F}^{2}+\|\bm{\gamma}^{(t)}-\bm{\gamma}^{(t-1)}\|^{2}+\|\bm{\psi}^{(t)}-\bm{\psi}^{(t-1)}\|^{2}\right)^{1/2}/\sqrt{J(G+2)+G(G-1)/2},$
[…]
$\max_{j\in\{1,...,J\}}h(|\lambda_{j2}^{(t)}|,...,|\lambda_{j,G+1}^{(t)}|).$
When both criteria are below their pre-specified thresholds, $\delta_{1}$ and $\delta_{2}$, respectively, we stop the algorithm. Let $T$ be the last iteration number. Then the estimated bi-factor model structure is given by
^$\widehat{\mathcal{B}}_{g}=\{j:|\lambda_{j,g+1}^{(T)}|>\delta_{2}\}.$
By our choice of the stopping criterion, the resulting ^$\widehat{\mathcal{B}}_{g}$, $g=1,...,G$, gives a partition of all the variables, and thus, the bi-factor structure is satisfied. For simulation studies in Section 3, we choose $\delta_{1}=\delta_{2}=10^{-2}$. For real data analysis in Section 4, we choose $\delta_{1}=\delta_{2}=10^{-4}$ to get a more accurate and reliable result.
The optimisation problem 7 is non-convex and can get stuck in a local minimum. Thus, we recommend running the proposed algorithm multiple times with random starting points and choosing the solution with the smallest objective function value. The algorithm can also suffer from slow convergence, especially when the penalty term becomes large. When the algorithm does not converge within $T_{\max}$ iterations, we suggest using the estimated parameters at the $T_{\max}$th iteration as the initial parameters and restarting the optimisation until a good proportion of them converge. In the simulation study in Section 3 below, the estimated parameters obtained using 50 random starting points are close to the global minimum in most cases in the simulation study. For the real data example in Section 4, 200 random starting points are used to ensure a reliable result. We set $T_{\max}=1000$ in all of our numerical studies.
2.3 Selecting the Number of Group Factors
In Sections 2.1 and 2.2, the number of group factors $G$ is treated as known. In practice, we can select its value based on the BIC (Schwarz,, 1978). Let $l_{G}$ denote the minimum loss function value in (2) when the number of group factors is $G$. As $l_{G}$ differs from twice the negative log-likelihood of the bi-factor model with $G$ group factors by a constant, and the numbers of nonzero parameters in $\Lambda$ and $\Psi$ do not depend on $G$, it is not difficult to see that the BIC of the bi-factor model with $G$ group factors differs from $l_{G}+({(G-1)G}\log(N))/2$ by a constant. Note that $(G-1)G/2$ is the number of free parameters in the correlation matrix $\Phi$. Thus, we write
$\text{BIC}_{G}=l_{G}+({(G-1)G}\log(N))/2.$
In practice, we choose the number of group factors $G$ from a candidate set $\mathcal{G}$. For each value of $G\in\mathcal{G}$, we run the ALM described in Section 2.2 to obtain the value of $l_{G}$. We then compute $\text{BIC}_{G}$ and choose $\widehat{G}$ with the smallest BIC value, i.e.,
$\widehat{G}=\operatornamewithlimits{arg\,min}_{G\in\mathcal{G}}~{}\text{BIC}_{G}.$
3 Simulation Study
3.1 Study I
In this study, we compare the proposed method with the oblique bi-factor rotation in Jennrich and Bentler, (2012) regarding their performance in the estimation accuracy of parameters and the recovery of the bi-factor structure. The oblique bi-factor rotation method first estimates the loading matrix $\widehat{\Lambda}$ under the exploratory factor analysis setting by the optimisation problem
$\widehat{\Lambda},\widehat{\psi}=\operatornamewithlimits{arg\,min}_{\Lambda\in\mathbb{R}^{J\times(G+1)},\psi\in\mathbb{R}^{J}}l(\Lambda\Lambda^{\top}+\Psi(\psi);S).$ (8)
We restrict $\lambda_{ij}=0$ for $i=2,\ldots,G$ and $j=i+1,\ldots,G+1$ to avoid the rotational indeterminacy of $\widehat{\Lambda}$, as suggested in Anderson and Rubin, (1956). Then the rotated solutions $\widehat{\Lambda}^{oblq}$ and $\widehat{\Phi}^{oblq}$ are obtained by finding a rotation matrix that solves the optimisation problem for oblique bi-factor rotation (Jennrich and Bentler,, 2012). The implementation in the R package GPArotation (Bernaards and Jennrich,, 2005) is used for solving this optimisation problem, which is based on a gradient projection algorithm. The optimisation problem for rotation is also nonconvex and thus may converge to local solutions. For a fair comparison, we also use 50 random starting points for the initial rotation matrix, as the same number of random starting points are used when running Algorithm 1.
We first examine the accuracy in estimating the loading matrix. We calculate the mean squared error (MSE) for $\widehat{\Lambda}$, after adjusting for the label and sign indeterminacy as considered in Theorem 1 and further discussed in Remark 1. More specifically, let $\mathcal{P}$ and $\mathcal{D}$ be the sets of permutation and sign flip matrices, respectively, as defined in Theorem 1. We define the MSE for $\Lambda$ as
$\text{MSE}_{\Lambda}=\min_{P\in\mathcal{P},D\in\mathcal{D}}\|\widehat{\Lambda}-\Lambda^{*}PD\|_{F}^{2}/(JK).$
This MSE can be applied to the loading matrix estimates from both methods.
To compare the two methods in terms of their performance in recovering the bi-factor structure, we derive a sparse loading structure from the rotated solution by hard thresholding, a procedure also performed in Jennrich and Bentler, (2012) for examining structure recovery. We let
^$\widehat{\mathcal{B}}_{g}^{oblq}=\{j:|\lambda_{j,g+1}^{oblq}|>\delta\},$
for $g=1,\ldots,G$ and some hard thresholding parameter $\delta>0$. We note that ^$\widehat{\mathcal{B}}_{g}^{oblq}$s may not yield an exact bi-factor structure in the sense that the current construction does not ensure them to be mutually exclusive.
Let $\{\mathcal{B}_{g}^{*},g=1,\ldots,G\}$ be the true non-overlapping group partitions of the $J$ variables, and let $\{\mathcal{B}_{g},g=1,\ldots,G\}$ be their estimates, either from the proposed method or the rotation method. As the group factors can only be recovered up to label swapping, as Theorem 1 suggests, we measure the matching between the true and estimated structure up to a permutation of the factor labels. Specifically, the following two evaluation criteria are considered:
  • Exact Match Criterion (EMC): $\max_{\sigma\in\widetilde{\mathcal{P}}}\prod_{g=1}^{G}\mathbf{1}(\mathcal{B}_{\sigma(g)}=\mathcal{B}_{g}^{*})$, which equals 1 when the bi-factor structure is correctly learned and 0 otherwise. Here, $(\sigma(1),...,\sigma(G))$ is a permutation of $1,...,G$, and $\widetilde{\mathcal{P}}$ is the set of all such permutations.
  • Average Correctness Criterion (ACC): $\max_{\sigma\in\widetilde{\mathcal{P}}}\sum_{g=1}^{G}(|\mathcal{B}_{g}^{*}\cap\mathcal{B}_{\sigma(g)}|+|\mathcal{B}_{\sigma(g)}^{C}\cap\mathcal{B}_{g}^{*C}|)/(JG)$, which is the average of the number of the bi-factor locations that are correctly learned. Here for any set $\mathcal{B}$, $\mathcal{B}^{C}=\{1,\ldots,J\}\setminus\mathcal{B}$ is the complement of set $\mathcal{B}$.
To examine the recovery of the bi-factor structure, for $(J,G)\in\{(15,3),(30,5)\}$, $N\in\{500,2000\}$ and $\delta\in\{0.2,0.4,0.6,0.8,1\}$, we consider 8 simulation settings. In the data generation model, $\mathcal{B}_{g}^{*}=\{g,g+G,\ldots,g+G(J/G-1)\}$ for $g=1,\ldots,G$, $\Psi^{*}=\mathbb{I}_{J\times J}$, $\Phi^{*}=\Phi^{*}(\bm{\gamma}^{*})$ follows the reparameterization in Section 2.2 with $\bm{\gamma}^{*}$ following i.i.d $N(0,1/4)$ and $\Lambda^{*}$ follows
[…]
for $j=1,\ldots,J$ and $k=1,\ldots,G+1$. Here $u_{jk}$ follows i.i.d Uniform(0,1) distribution and $x_{jk}$ follows i.i.d standard normal distribution.
For each setting, we first generate $\Lambda^{*}$ and $\Phi^{*}$ once and use them to generate 100 datasets. The result on the estimation of the loading matrix is shown in Figure 1, and the results of the recovery of the bi-factor structure are shown in Tables 1 and 2. It shows that the proposed method outperforms the rotation method in terms of both criteria. Specifically, according to the EMC criterion, the proposed method has a good chance to completely recover the bi-factor structure, especially when $G$ is small and $N$ is large, while the rotation method with hard thresholding has a much lower chance of achieving so. Regarding the ACC criterion, it suggests that both methods can identify the group factors reasonably well, while the proposed method is more accurate.
Figure 1: Simulation results of the MSE for $\widehat{\Lambda}$ of ALM and exploratory bi-factor rotation.
Table 1: Simulation results of the EMC of ALM and exploratory bi-factor rotation.
$(J,G)$ $N$ ALM $\delta=0.2$ $\delta=0.4$ $\delta=0.6$ $\delta=0.8$ $\delta=1$
(15,3) 500 0.9 0.36 0.00 0.00 0.00 0.00
2000 1.00 0.62 0.00 0.00 0.00 0.00
(30,5) 500 0.70 0.00 0.00 0.00 0.00 0.00
2000 0.96 0.00 0.00 0.00 0.00 0.00
Table 2: Simulation results of the ACC of ALM and exploratory bi-factor rotation.
$(J,G)$ $N$ ALM $\delta=0.2$ $\delta=0.4$ $\delta=0.6$ $\delta=0.8$ $\delta=1$
(15,3) 500 0.99 0.98 0.95 0.93 0.90 0.87
2000 1.00 0.99 0.96 0.93 0.91 0.89
(30,5) 500 0.99 0.84 0.95 0.97 0.95 0.94
2000 0.99 0.89 0.97 0.97 0.95 0.94
3.2 Study II
In this study, we examine the selection of the number of factors by $\text{BIC}_{G}$ in Section 2.3. We compare it with selecting the number of factors under the exploratory factor analysis model without assuming a bi-factor structure. For the proposed method, we set the candidate set $\mathcal{G}=\{G^{*}-1,G^{*},G^{*}+1\}$, where $G^{*}$ is the true number of group factors. For exploratory factor analysis, we also use the BIC for determining the number of factors, which is defined as
$\text{BIC}_{K}^{e}=l_{K}^{e}+(JK-K(K-1)/2)\log(N),$
where $K$ is the number of factors in the exploratory factor analysis model, and $l_{K}^{e}=l(\widehat{\Lambda}\widehat{\Lambda}^{\top}+\Psi(\widehat{\psi});S)$ with $\widehat{\Lambda}$ and $\widehat{\psi}$ from (8). As the number of factors in the exploratory factor analysis model equals the number of group factors plus one, we choose $K$ from the candidate set $\mathcal{K}=\{G+1:G\in\mathcal{G}\}$. Let $\widehat{K}=\operatornamewithlimits{arg\,min}_{K\in\mathcal{K}}~{}\text{BIC}_{K}^{e}.$ Then the estimate of $G$ by exploratory factor analysis is $\widehat{G}=\widehat{K}-1$. The selection accuracy is evaluated by the selection correctness (SC) criterion, defined as $\mathbf{1}(\widehat{G}=G^{*})$, where $\widehat{G}$ is obtained using the proposed method in Section 2.3 or under the exploratory factor analysis model described above.
We conduct simulations under four settings, with $(J,G^{*})\in\{(15,3),(30,5)\}$ and $N\in\{500,2000\}$. The same data generation model in Study I is used. Under each setting, 100 independent simulations are performed. The results are given in Table 3, where the column indicated by $\bar{G}$ reports the average value of $\widehat{G}$. We see that both methods can select the number of factors reasonably well, with their accuracy being 100% when $G^{*}=3$ for both sample sizes and when $G^{*}=5$ for the larger sample size $N=2000$. This is not surprising, given that the BIC has asymptotic consistency in selecting the number of factors under both models. When $G^{*}=5$ and for the smaller sample size $N=500$, which is the most challenging setting, the proposed method achieves an accuracy of 93%, while that of the exploratory factor analysis method is only 81%. It suggests that the proposed method has a better performance in smaller sample settings. This result is expected, as the exploratory factor analysis method doesn’t utilize the information about the bi-factor structure. Consequently, it overestimates the number of parameters, which leads to a larger penalty term and, subsequently, a tendency to under-select $G$.
Table 3: Simulation results of the selection of number of factors by BIC.
ALM Exploratory
$(J,G^{*})$ $N$ $\bar{G}$ SC $\bar{G}$ SC
(15,3) 500 3 1 3 1
2000 3 1 3 1
(30,5) 500 5.07 0.93 4.81 0.81
2000 5 1 5 1
4 Real Data Analysis
In this section, we apply the exact exploratory bi-factor analysis to a personality assessment dataset based on the International Personality Item Pool (IPIP) NEO 120 personality inventory (Johnson,, 2014) [^1] The data are downloaded from https://osf.io/tbmh5/. We investigate the structure of the Extraversion scale based on a sample of 1,107 UK male participants aged between 25 and 30 years. This scale consists of 24 items, which are designed to measure six facets of Extraversion, including Friendliness (E1), Gregariousness (E2), Assertiveness (E3), Activity Level(E4), Excitement-Seeking (E5) and Cheerfulness (E6); see Section B for the details. All the items are on a 1-5 Likert scale, and the reversely worded items have been reversely scored so that a larger score always means a higher level of extraversion. There is no missing data. Detailed descriptions of the items can be found in the Appendix B.
Using a candidate set $\mathcal{G}=\{2,\ldots,12\}$, the BIC procedure given in Section 2.3 selects seven group factors. The estimated loading matrix is given in Table 4, and the estimated factor correlation matrix is given below. We point out that the estimated bi-factor structure does not meet Condition 3, one of the sufficient conditions for Theorem 1. However, as shown in the Appendix C.2, with an additional mild assumption, this structure and its parameters are still identifiable.
We now examine the estimated model. We first notice that the loadings on the general factor are all positive. Consequently, this factor can be easily interpreted as the general extraversion factor. The seven group factors are closely related to the six aspects of extraversion. Specifically, we interpret the group factors G1, G3, G4 and G5 as the Friendliness, Cheerfulnes, Assertiveness, and Excitement-Seeking factors, respectively, as the items loading on them highly overlap with the items that are used to define the corresponding aspects. In particular, the items loading on G3 and G5 are exactly those that define the Cheerfulness and Excitement-Seeking aspects, respectively. The items loading on G1 include all the items that define the Friendliness aspect and an additional item “7. Prefer to be alone”, a negatively worded item that is used to define the Gregariousness aspect. This additional item aligns well with the Friendliness dimension, given the social nature behind it. In addition, the items loading on G4 consist of all the items that define the Assertiveness aspect and an additional item “6. Talk to a lot of different people at parties”, which is used to define the Gregariousness aspect. This additional item aligns with the Assertiveness dimension in that talking to many different people at parties typically requires sufficient confidence, a key element of Assertiveness.
Table 4: Estimated bi-factor loading matrix $\widehat{\Lambda}$ with seven group factors.
Items Sign General G1 G2 G3 G4 G5 G6 G7
1 +E1 0.85 0.26 0 0 0 0 0 0
2 +E1 0.73 0.48 0 0 0 0 0 0
3 --E1 0.74 0.57 0 0 0 0 0 0
4 --E1 0.68 0.58 0 0 0 0 0 0
5 +E2 0.94 0 0 0 0 0 0.26 0
6 +E2 1.01 0 0 0 0.17 0 0 0
7 --E2 0.53 0.52 0 0 0 0 0 0
8 --E2 0.67 0 0 0 0 0 1.06 0
9 +E3 0.37 0 0 0 0.86 0 0 0
10 +E3 0.38 0 0 0 0.81 0 0 0
11 +E3 0.28 0 0 0 0.74 0 0 0
12 --E3 0.39 0 0 0 0.75 0 0 0
13 +E4 0.20 0 0 0 0 0.81 0 0
14 +E4 0.40 0 0 0 0 0.82 0 0
15 +E4 0.40 0 0 0 0 0.60 0 0
16 --E4 0.04 0 0 0 0 0.47 0 0
17 +E5 0.46 0 0 0 0 0 0 0.46
18 +E5 0.47 0 0 0 0 0 0 0.71
19 +E5 0.35 0 0.86 0 0 0 0 0
20 +E5 0.56 0 0.71 0 0 0 0 0
21 +E6 0.59 0 0 0.42 0 0 0 0
22 +E6 0.64 0 0 0.48 0 0 0 0
23 +E6 0.46 0 0 0.76 0 0 0 0
24 +E6 0.41 0 0 0.74 0 0 0 0
$\widehat{\Phi}=\left(\begin{matrix}1&0&0&0&0&0&0&0\\ 0&1&-0.24&0.54&0.37&0.16&0.51&0.08\\ 0&-0.24&1&-0.04&0.05&-0.09&-0.01&0.51\\ 0&0.55&-0.04&1&0.30&0.28&0.20&0.25\\ 0&0.37&0.05&0.30&1&0.38&0.15&0.29\\ 0&0.16&-0.09&0.28&0.38&1&0.11&0.22\\ 0&0.51&-0.01&0.20&0.15&0.11&1&0.10\\ 0&0.08&0.51&0.25&0.29&0.22&0.10&1\\ \end{matrix}\right).$ (10)
The group factors G2 and G7 may be viewed as two subdimensions of the Excitement-Seeking aspect, as each of them is loaded with two items that define the Excitement-Seeking aspect. Specifically, G2 is loaded with the items “19. Enjoy being reckless” and “20. Act wild and crazy”, while G7 is loaded with the items “17. Love excitement” and “18. Seek adventure”. We believe that G2 emphasises the thrill of the moment of excitement and the disregard for potential consequences, while G7 emphasizes the pursuit of meaningful and fulfilling experiences. Therefore, we interpret G2 as the Reckless Excitement-Seeking factor, while interpret G7 as the Meaningful Excitement-Seeking factor. Finally, G6 is loaded with two items, “5. Love large parties” and “8. Avoid crowds”, where item 8 is reversely worded. Both items are used to define the Gregariousness aspect. Compared with items 6 and 7, which are also used to define the Gregariousness aspect but now load on two different group factors, these two items may better reflect the essence of Gregariousness – the tendency to enjoy the company of others. We thus interpret G6 as the Gregariousness factor. We also notice that most correlations between the group factors are positive, except for some of the correlations with G2. Specifically, we see that G2 (Reckless Excitement-Seeking) has a moderate negative correlation with G1 (Friendliness) while a reasonably high correlation with G7 (Meaningful Excitement-Seeking).
5 Discussions
This paper proposes a constraint-based optimisation method for exploratory bi-factor analysis. This method turns the problem of exploratory bi-factor analysis into an equality-constrained optimisation problem in a continuous domain and solves this optimisation problem by an augmented Lagrangian method. Compared with the rotation method of Jennrich and Bentler, (2011, 2012), the proposed method can learn an exact loading structure without a post-hoc treatment step. In the simulation studies, the ALM method achieves higher estimation accuracy and has a greater likelihood of recovering the entire bi-factor structure compared to the previous rotation method. Additionally, the ALM method correctly estimates the number of the group factors in most of the simulation replications. In the real data analysis concerning the Extraversion scale, the ALM method explores a bi-factor structure with seven group factors. Moreover, the selected group factors are psychologically interpretable, demonstrating the utility of our proposed method.
An innovation of current research is turning a model selection problem, which is combinatory by nature, into a continuous optimization problem. This avoids a computationally intensive search procedure for fitting many possible models and comparing their fits, noting that the number of possible models grows exponentially with $J$. We admit that this continuous optimization formulation also has a limitation. The space for the bi-factor loading matrix characterised by the nonlinear equality constraints in (1) is highly nonconvex, and thus, the ALM may sometimes converge to a local minimum. To alleviate this issue, we suggest running the ALM with multiple random starting points and then choosing the solution with the smallest objective function value. Based on our simulation results, using 50 starting points seems sufficient to converge to somewhere close to the true parameters up to a label swapping of the group factors and a sign indeterminacy of loadings in almost all replications under the settings considered in the simulation study.
This research leads to several new directions for exploratory analysis of factor models with structure constraints on the loading matrix. First, as pointed out earlier, the proposed approach can be easily adapted to non-linear bi-factor models for dichotomous, ordinal, and nominal data. Under the confirmatory setting, these models are typically estimated by maximising the marginal log-likelihood function or other objective functions (e.g., a composite likelihood). Under the exploratory setting, one only needs to maximise the same objective function subject to the same bi-factor constraints in (1), for which the ALM adapts naturally. It is worth noting that, however, the marginal likelihood of the non-linear bi-factor models typically involves multidimensional integrals with respect to the factors, and they do not have an analytic form. Consequently, solving the Lagrangian augmented objective functions using the standard expectation-maximisation (EM) algorithm (Dempster et al.,, 1977; Bock and Aitkin,, 1981) can be computationally intensive. One possible solution is to use a stochastic approximation method (Zhang and Chen,, 2022; Oka et al.,, 2024). These methods avoid the high computational cost of numerical integrals in the expectation-maximisation algorithm by constructing stochastic gradients of the marginal log-likelihood through Markov chain Monte Carlo sampling.
Second, the proposed constraints can also be combined with exploratory factor analysis techniques to learn a bi-factor structure in two steps. Suppose an initial loading matrix estimate $\widehat{\Lambda}$ has been obtained under the constraint that the factors are orthogonal (i.e., $\Phi$ is an identity matrix). It may be obtained by a standard exploratory factor analysis method. In that case, we can find a bi-factor structure that best approximates $\widehat{\Lambda}$ (up to a rotation) by minimising $\|\Lambda\Phi(\bm{\gamma})\Lambda^{\top}-\widehat{\Lambda}(\widehat{\Lambda})^{\top}\|_{F}$ with respect to $\Lambda$ and $\bm{\gamma}$ under the constraints in (1). This optimisation can again be solved by an ALM.
Finally, as we demonstrate in Appendix A, the set of constraints in (1) can be extended to characterise the loading structure of a hierarchical factor model (Schmid and Leiman,, 1957; Yung et al.,, 1999), which can be used to learn a hierarchical factor structure. This exploratory hierarchical factor analysis may allow researchers to learn more refined and interpretable latent structures from psychometric data. However, one should note that exploratory hierarchical factor analysis is more complex than exploratory bi-factor analysis, as the factor hierarchy in the former can be much more complex than the two-layer hierarchy in the latter. The learning algorithm in Appendix A requires the factor hierarchy to be known (see, e.g., Panel (b) of Figure A.1). When it is unknown, one may learn the hierarchical factor structure, including the factor hierarchy as a part, by combining an iterative procedure that adaptively adds hierarchical factors into the model and a true pruning method for removing the redundant factors. Adding hierarchical factors can be done by using the learning algorithm discussed in Appendix A, and tree pruning may be achieved by a model selection procedure based on the BIC. We leave this problem for future investigation.
Appendix
Appendix A Extension to Hierarchical Factor Analysis
A.1 Constrained Optimisation for Exploratory Hierarchical Factor Analysis
To further demonstrate the advantages of the constraint-based approach, we discuss how it can be extended for exploratory hierarchical factor analysis. Following the terminology adopted in Yung et al., (1999), we consider general hierarchical factor models. Such a model has several layers of factors. In each layer, each observed variable loads on exactly one of the factors in that layer. The numbering of the layers is determined by the number of factors in the layer, starting from the layer with the largest number of factors. Each factor in a lower layer is nested within a factor in a higher layer, in the sense that the variables loading on the lower-layer factor must also all load on a higher-layer factor. All the factors are assumed to be uncorrelated (i.e., $\Phi$ is an identity matrix), though this assumption may be relaxed to allow some correlations between factors within the same layer as in the extended bi-factor model.
[…]
(a) The path diagram of a three-layer hierarchical factor model.
[…]
(b) The corresponding factor hierarchy.
Figure A.1: The illustrative example of a three-layer hierarchical factor model.
Panel (a) of Figure A.1 provides the path diagram of a hierarchical factor model that has three layers, with factor $F_{1}$ in layer 3, factors $F_{2}$ and $F_{3}$ in layer 2, and factors $F_{4}$-$F_{7}$ in layer 1. The corresponding factor hierarchy is summarised in Panel (b) of Figure A.1 that takes the form of a tree, where $F_{2}$ and $F_{3}$ are nested within $F_{1}$, $F_{4}$ and $F_{5}$ are nested within $F_{2}$, and $F_{6}$ and $F_{7}$ are nested within $F_{3}$. In what follows, we show how the loading structure of this three-layer hierarchical model can be learned by a constrained optimisation method, assuming that the factor hierarchy in Panel (b) of Figure A.1 is known while the variables loading on each factor are unknown. The goal is to learn how the observed variables load on the seven factors.
Following the same notation for bi-factor analysis, the population covariance matrix of observed variables under the hierarchical factor model can be written as
$\Sigma=\Lambda\Lambda^{\top}+\Psi,$
where $\Lambda$ is a $J\times 7$ matrix, and $\Psi$ is a $J\times J$ diagonal matrix. Note that we no longer need the correlation matrix $\Phi$ in the expression as it is now an identity matrix. The constraints implied by the hierarchical factor structure become:
$\lambda_{j2}\lambda_{j3}=0,~{}~{}\lambda_{j2}\lambda_{j6}=0,~{}~{}\lambda_{j2}\lambda_{j7}=0,$ (A.1)
$\lambda_{j3}\lambda_{j4}=0,~{}~{}\lambda_{j3}\lambda_{j5}=0,$
$\lambda_{j4}\lambda_{j5}=0,~{}~{}\lambda_{j6}\lambda_{j7}=0,~{}~{}j=1,...,J.$
Consequently, the corresponding hierarchical factor model can be learned by minimising the loss function $l(\Lambda\Lambda^{\top}+\Psi(\bm{\psi});S)$, subject to the constraints in (A.1).
Although the above discussion focuses on the specific hierarchical factor structure in Figure A.1, when given a different factor hierarchy, it is easy to derive similar constraints as in (A.1) by induction. Based on the constraints, the corresponding hierarchical factor model can be learned by an ALM.
Finally, we note that the factor hierarchy is typically unknown in practice. In that case, we need an algorithm that simultaneously learns the factor hierarchy and the variable loadings on the hierarchical factors. As there are exponentially many choices for the structure of factor hierarchy, this problem is more challenging than the setting when the factor hierarchy is known. It is also more challenging than exploratory bi-factor analysis with unknown group factors, as the bi-factor model has a simple two-layer factor hierarchy that is completely determined by the number of factors. See further discussions on this learning problem in Section 5.
A.2 Simulation
In this section, we examine the recovery of the hierarchical structure of our method. For $J\in\{20,40\}$ and $N\in\{500,2000\}$, a data generation model is considered, resulting in a total 3 simulation settings. With slight abuse of notation, we denote by $\mathcal{B}_{g}^{*}$ as the true item groups related with the $g$th factor. In the data generation model, $B_{1}^{*}=\{1,\ldots,J\}$, $B_{2}^{*}=\{1,\ldots,J/2\}$, $B_{3}^{*}=\{J/2,\ldots,J\}$, $B_{4}^{*}=\{1,\ldots,J/4\}$, $B_{5}^{*}=\{J/4,\ldots,J/2\}$, $B_{6}^{*}=\{J/2,\ldots,3J/4\}$, $B_{7}^{*}=\{3J/4,\ldots,J\}$. $\Psi^{*}=\mathbb{I}_{J\times J}$, and $\Lambda^{*}$ follows
[…]
for $j=1,\ldots,J$ and $k=1,\ldots,G+1$. Here $u_{jk}$ follows i.i.d Uniform(0,1) distribution and $x_{jk}$ follows i.i.d standard normal distribution.
The estimated parameters $\widehat{\Lambda}$ and $\widehat{\Psi}$ follow the same ALM algorithm in Section 2.2 except that the distance between the estimate and the space of the hierarchical factor loading matrices measured by
$\max_{j\in\{1,...,J\}}\widetilde{h}(|\lambda_{j2}^{(t)}|,...,|\lambda_{j,G+1}^{(t)}|),$
where the function $\widetilde{h}$ returns the third-largest value of a vector. The estimated hierarchical factor model structure is then given by
^$\widehat{\mathcal{B}}_{g}=\{j:|\lambda_{j,g+1}^{(T)}|>\delta_{2}\}.$
We also choose $\delta_{2}=10^{-2}$ on the following simulation study.
Since label-switching problem exists in factors that are nested within the same hierarchical factor, there exists 8 possible permutations of labels resulting in the same hierarchical structure. We denote by $\mathcal{R}$ as the set of the 8 permutations. Then, the evaluation criteria for the recovery of the hierarchical structure are defined as:
  • Exact Match Criterion(EMC): $\max_{\sigma\in\mathcal{R}}\prod_{g=1}^{G}\mathbf{1}(\mathcal{B}_{\sigma(g)}=\mathcal{B}_{g}^{*})$, which equals 1 when the bi-factor structure is correctly learned and 0 otherwise.
  • Average Correctness Criterion(ACC): $\max_{\sigma\in\mathcal{R}}\sum_{g=1}^{G}(|\mathcal{B}_{g}^{*}\cap\mathcal{B}_{\sigma(g)}|+|\mathcal{B}_{\sigma(g)}^{C}\cap\mathcal{B}_{g}^{*C}|)/(JG)$.
For each setting, we first generate $\Lambda^{*}$ once and use them to generate 100 datasets. The averaged results under 100 replication are shown in Table A.1. From the simulation results, we find that our method performs well on the recovery of hierarchical structure.
Table A.1: Simulation results of the recovery of hierarchical factor structure.
$J$ $N$ EMC ACC
20 500 0.83 0.99
2000 0.92 0.99
40 500 1.00 1.00
2000 1.00 1.00
Appendix B Extraversion Scale Item Key
Table B.1: Extraversion Item Key
Item Sign Facet Item
1 +E1 Friendliness Make friends easily.
2 +E1 Friendliness Feel comfortable around people.
3 --E1 Friendliness Avoid contacts with others.
4 --E1 Friendliness Keep others at a distance.
5 +E2 Gregariousness Love large parties.
6 +E2 Gregariousness Talk to a lot of different people at parties.
7 --E2 Gregariousness Prefer to be alone.
8 --E2 Gregariousness Avoid crowds.
9 +E3 Assertiveness Take charge.
10 +E3 Assertiveness Try to lead others.
11 +E3 Assertiveness Take control of things.
12 --E3 Assertiveness Wait for others to lead the way.
13 +E4 Activity Level Am always busy.
14 +E4 Activity Level Am always on the go.
15 +E4 Activity Level Do a lot in my spare time.
16 --E4 Activity Level Like to take it easy.
17 +E5 Excitement-Seeking Love excitement.
18 +E5 Excitement-Seeking Seek adventure.
19 +E5 Excitement-Seeking Enjoy being reckless.
20 +E5 Excitement-Seeking Act wild and crazy.
21 +E6 Cheerfulness Radiate joy.
22 +E6 Cheerfulness Have a lot of fun.
23 +E6 Cheerfulness Love life.
24 +E6 Cheerfulness Look at the bright side of life.
Appendix C Technical Proofs
C.1 Proof of Theorem 1
Suppose that $\Lambda\Phi(\Lambda)^{\top}+\Psi=\Lambda^{*}\Phi^{*}(\Lambda^{*})^{\top}+\Psi^{*}$. Under Condition 1, we have $\Lambda\Phi(\Lambda)^{\top}=\Lambda^{*}\Phi^{*}(\Lambda^{*})^{\top}$. For the simplicity of the notation, we substitute $\Lambda[\mathcal{B}^{*}_{g},\{1,\ldots,G+1\}]$ for $\Lambda[\mathcal{B}^{*}_{g},:]$. The proof consists of three parts:(1) show the bi-factor structure of $\mathcal{B}_{g_{1}}^{*}$ is unique, (2) show that combined with some group $g_{2}\in\mathcal{H}^{*},g_{2}\neq g_{1}$, $\Lambda[\mathcal{B}^{*}_{g_{1}},:]$ is identified up to a sign flip and a group permutation, (3) complete the proof of Theorem 1.
We first consider the equation
$\Lambda[\mathcal{B}^{*}_{g_{1}},:]\Phi(\Lambda[\mathcal{B}^{*}_{g_{1}},:])^{\top}=\Lambda^{*}[\mathcal{B}^{*}_{g_{1}},\{1,1+g_{1}\}](\Lambda^{*}[\mathcal{B}^{*}_{g_{1}},\{1,1+g_{1}\}])^{\top}.$ (C.1)
Since the right side matrix has rank 2, there exist 2 possible bi-factor structures of the left side: ($\Lambda[\mathcal{B}^{*}_{g_{1}},\{1,1+g^{\prime}_{1}\}](\Lambda[\mathcal{B}^{*}_{g_{1}},\{1,1+g^{\prime}_{1}\}])^{\top}=\Lambda^{*}[\mathcal{B}^{*}_{g_{1}},\{1,1+g_{1}\}](\Lambda^{*}[\mathcal{B}^{*}_{g_{1}},\{1,1+g_{1}\}])^{\top}$ for some $g^{\prime}_{1}\in\{1,\ldots,G\}$and (2) There exists a partition of $\mathcal{B}^{*}_{g_{1}}=\mathcal{B}^{*}_{g_{1,1}}\cup\mathcal{B}^{*}_{g_{1,2}}$ and $g^{\prime}_{1},g^{\prime}_{2}\in\{1,\ldots,G\}$ such that
$\begin{pmatrix}\bm{\lambda}_{1}&\bm{\lambda}_{g^{\prime}_{1}}&\mathbf{0}\\ \bm{\lambda}_{2}&\mathbf{0}&\bm{\lambda}_{g^{\prime}_{2}}\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&\phi_{1+g^{\prime}_{1},1+g^{\prime}_{2}}\\ 0&\phi_{1+g^{\prime}_{1},1+g^{\prime}_{2}}&1\end{pmatrix}\begin{pmatrix}\bm{\lambda}_{1}^{\top}&\bm{\lambda}_{2}^{\top}\\ \bm{\lambda}_{g^{\prime}_{1}}^{\top}&\mathbf{0}^{\top}\\ \mathbf{0}^{\top}&\bm{\lambda}_{g^{\prime}_{2}}^{\top}\end{pmatrix}=\begin{pmatrix}\bm{\lambda}^{*}_{1}&\bm{\lambda}^{*}_{g_{1}}\\ \bm{\lambda}^{*}_{2}&\bm{\lambda}^{*}_{g_{2}}\end{pmatrix}\begin{pmatrix}(\bm{\lambda}^{*}_{1})^{\top}&(\bm{\lambda}^{*}_{2})^{\top}\\ (\bm{\lambda}^{*}_{g_{1}})^{\top}&(\bm{\lambda}^{*}_{g_{2}})^{\top}\end{pmatrix},$ (C.2)
where $\bm{\lambda}_{i}=\Lambda[\mathcal{B}^{*}_{g_{1,i}},\{1\}]$, $\bm{\lambda}_{g^{\prime}_{i}}=\Lambda[\mathcal{B}^{*}_{g_{1,i}},\{1+g^{\prime}_{i}\}]$, $\bm{\lambda}^{*}_{i}=\Lambda^{*}[\mathcal{B}^{*}_{g_{1,i}},\{1\}]$ and $\bm{\lambda}^{*}_{g_{i}}=\Lambda^{*}[\mathcal{B}^{*}_{g_{1,i}},\{1+g_{1}\}]$ with $\bm{\lambda}_{g^{\prime}_{i}}\neq\mathbf{0}$ for $i=1,2$.
Here we consider the second case. Since the matrix on the right side has rank 2, we must have $(\bm{\lambda}_{i},\bm{\lambda}_{g^{\prime}_{i}})$ has rank 1 for $i=1,2$, which leads to the fact that $(\bm{\lambda}^{*}_{i},\bm{\lambda}^{*}_{g_{i}})$ has rank 1. However, by Condition 2, there exists at least one of $(\bm{\lambda}^{*}_{1},\bm{\lambda}^{*}_{g_{1}})$ and $(\bm{\lambda}^{*}_{2},\bm{\lambda}^{*}_{g_{2}})$ has rank 2. Thus, we must have $\Lambda[\mathcal{B}^{*}_{g_{1}},\{1,1+g^{\prime}_{1}\}](\Lambda[\mathcal{B}^{*}_{g_{1}},\{1,1+g^{\prime}_{1}\}])^{\top}=\Lambda^{*}[\mathcal{B}^{*}_{g_{1}},\{1,1+g_{1}\}](\Lambda^{*}[\mathcal{B}^{*}_{g_{1}},\{1,1+g_{1}\}])^{\top}$ for some $g^{\prime}_{1}\in\{1,\ldots,G\}$. Without loss of generation, we assume $g^{\prime}_{1}=g_{1}$.
Secondly, there exits some $g_{2}\in\mathcal{H}^{*}$ and $g_{2}\neq g_{1}$ by Condition 2. We consider the $\mathcal{B}^{*}_{g_{1}}\cup\mathcal{B}^{*}_{g_{2}}$ rows and $\mathcal{B}^{*}_{g_{1}}\cup\mathcal{B}^{*}_{g_{2}}$ columns of $\Lambda\Phi(\Lambda)^{\top}$ and $\Lambda^{*}\Phi^{*}(\Lambda^{*})^{\top}$. Since the bi-factor structure of the $\mathcal{B}^{*}_{g_{1}}$ rows and $\mathcal{B}^{*}_{g_{1}}$ columns has already been known, there are two possible bi-factor structures: (1) There exists some $g^{\prime}_{2}\in\{1,\ldots,G\}$ such that
$\begin{pmatrix}\bm{\lambda}_{1}&\bm{\lambda}_{g_{1}}&\mathbf{0}\\ \bm{\lambda}_{2}&\mathbf{0}&\bm{\lambda}_{g^{\prime}_{2}}\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&\rho_{1,2}\\ 0&\rho_{1,2}&1\end{pmatrix}\begin{pmatrix}\bm{\lambda}_{1}^{\top}&\bm{\lambda}_{2}^{\top}\\ \bm{\lambda}_{g_{1}}^{\top}&\mathbf{0}^{\top}\\ \mathbf{0}^{\top}&\bm{\lambda}_{g^{\prime}_{2}}^{\top}\end{pmatrix}=\begin{pmatrix}\bm{\lambda}^{*}_{1}&\bm{\lambda}^{*}_{g_{1}}&\mathbf{0}\\ \bm{\lambda}^{*}_{2}&\mathbf{0}&\bm{\lambda}^{*}_{g_{2}}\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&\rho_{1,2}^{*}\\ 0&\rho_{1,2}^{*}&1\end{pmatrix}\begin{pmatrix}(\bm{\lambda}_{1}^{*})^{\top}&(\bm{\lambda}_{2}^{*})^{\top}\\ (\bm{\lambda}_{g_{1}}^{*})^{\top}&\mathbf{0}^{\top}\\ \mathbf{0}^{\top}&(\bm{\lambda}_{g_{2}}^{*})^{\top},\end{pmatrix}$ (C.3)
where $\bm{\lambda}_{i}=\Lambda[\mathcal{B}^{*}_{g_{i}},\{1\}]$, $\bm{\lambda}_{g^{\prime}_{i}}=\Lambda[\mathcal{B}^{*}_{g_{i}},\{1+g^{\prime}_{i}\}]$, $\bm{\lambda}^{*}_{i}=\Lambda^{*}[\mathcal{B}^{*}_{g_{i}},\{1\}]$ and $\bm{\lambda}^{*}_{g_{i}}=\Lambda^{*}[\mathcal{B}^{*}_{g_{i}},\{1+g_{1}\}]$ for $i=1,2$. $\rho_{1,2}=\phi_{1+g^{\prime}_{1},1+g^{\prime}_{2}}$ and $\rho^{*}_{1,2}=\phi^{*}_{1+g_{1},1+g_{2}}$.
(2) There exists a partition of $\mathcal{B}^{*}_{g_{2}}=\mathcal{B}^{*}_{g_{2,1}}\cup\mathcal{B}^{*}_{g_{2,2}}$ and $g^{\prime}_{2}\in\{1,\ldots,G\}$ such that
$\begin{pmatrix}\bm{\lambda}_{1}&\bm{\lambda}_{g_{1}}&\mathbf{0}\\ \bm{\lambda}_{2,1}&\bm{\lambda}_{2,g_{1}}&\mathbf{0}\\ \bm{\lambda}_{2,2}&\mathbf{0}&\bm{\lambda}_{2,g^{\prime}_{2}}\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&\rho_{1,2}\\ 0&\rho_{1,2}&1\end{pmatrix}\begin{pmatrix}\bm{\lambda}_{1}^{\top}&\bm{\lambda}_{2,1}^{\top}&\bm{\lambda}_{2,2}^{\top}\\ \bm{\lambda}_{g_{1}}^{\top}&\bm{\lambda}_{2,g_{1}}^{\top}&\mathbf{0}^{\top}\\ \mathbf{0}^{\top}&\mathbf{0}^{\top}&\bm{\lambda}_{2,g^{\prime}_{2}}^{\top}\end{pmatrix}$ (C.4)
$=\begin{pmatrix}\bm{\lambda}^{*}_{1}&\bm{\lambda}^{*}_{g_{1}}&\mathbf{0}\\ \bm{\lambda}^{*}_{2,1}&\mathbf{0}&\bm{\lambda}^{*}_{g_{2},1}\\ \bm{\lambda}^{*}_{2,2}&\mathbf{0}&\bm{\lambda}^{*}_{g_{2},2}\\ \end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&\rho_{1,2}^{*}\\ 0&\rho_{1,2}^{*}&1\end{pmatrix}\begin{pmatrix}(\bm{\lambda}_{1}^{*})^{\top}&(\bm{\lambda}_{2,1}^{*})^{\top}&(\bm{\lambda}_{2,2}^{*})^{\top}\\ (\bm{\lambda}_{g_{1}}^{*})^{\top}&\mathbf{0}^{\top}&\mathbf{0}^{\top}\\ \mathbf{0}^{\top}&(\bm{\lambda}_{g_{2},1}^{*})^{\top}&(\bm{\lambda}_{g_{2},2}^{*})^{\top}\end{pmatrix},$
where $\bm{\lambda}_{1}=\Lambda[\mathcal{B}^{*}_{g_{1}},\{1\}]$, $\bm{\lambda}_{g_{1}}=\Lambda[\mathcal{B}^{*}_{g_{1}},\{1+g_{1}\}]$, $\bm{\lambda}^{*}_{1}=\Lambda^{*}[\mathcal{B}^{*}_{g_{1}},\{1\}]$, $\bm{\lambda}^{*}_{g_{1}}=\Lambda^{*}[\mathcal{B}^{*}_{g_{1}},\{1+g_{1}\}]$, $\bm{\lambda}_{2,i}=\Lambda[\mathcal{B}^{*}_{g_{2},i},\{1\}]$, $\bm{\lambda}^{*}_{2,i}=\Lambda[\mathcal{B}^{*}_{g_{2},i},\{1\}]$, $\bm{\lambda}^{*}_{g_{2},i}=\Lambda[\mathcal{B}^{*}_{g_{2},i},\{1+g_{2}\}]$ for $i=1,2$, $\bm{\lambda}_{2,g_{1}}=\Lambda[\mathcal{B}^{*}_{g_{2},1},\{1+g_{1}\}]$, $\bm{\lambda}_{2,g^{\prime}_{2}}=\Lambda[\mathcal{B}^{*}_{g_{2},2},\{1+g^{\prime}_{2}\}]$, $\rho_{1,2}=\phi_{1+g^{\prime}_{1},1+g^{\prime}_{2}}$ and $\rho^{*}_{1,2}=\phi^{*}_{1+g^{\prime}_{1},1+g_{2}}$.
For the second case, there exists some $\alpha$ such that $\bm{\lambda}_{1}=\cos{\alpha}\bm{\lambda}^{*}_{1}-\sin{\alpha}\bm{\lambda}^{*}_{g_{1}}$ and $\bm{\lambda}_{g_{1}}=\sin{\alpha}\bm{\lambda}^{*}_{1}+\cos{\alpha}\bm{\lambda}^{*}_{g_{1}}$. Since the $\mathcal{B}_{g_{2}}^{*}$ rows and the $\mathcal{B}_{g_{2}}^{*}$ columns of $\Lambda\Phi(\Lambda)^{\top}$ and $\Lambda^{*}\Phi^{*}(\Lambda^{*})^{\top}$ have rank 2, under the bi-factor structure of the second case, we have $(\bm{\lambda}_{2,1},\bm{\lambda}_{2,g_{1}},\bm{\lambda}^{*}_{2,1},\bm{\lambda}^{*}_{g_{2},1})$ has rank 1 and $(\bm{\lambda}_{2,2},\bm{\lambda}_{2,g^{\prime}_{2}},\bm{\lambda}^{*}_{2,2},\bm{\lambda}^{*}_{g_{2},2})$ has rank 1 and $\bm{\lambda}^{*}_{g_{2},2}\neq\mathbf{0}$. We then assume that $\lambda^{*}_{2,2}=k_{1}\bm{\lambda}^{*}_{g_{2},2}$, $\bm{\lambda}_{2,2}=k_{2}\bm{\lambda}^{*}_{g_{2},2}$ and $\bm{\lambda}_{2,g^{\prime}_{2}}=k_{3}\bm{\lambda}^{*}_{g_{2},2}$. For the $\mathcal{B}^{*}_{g_{2},2}$ rows and the $\mathcal{B}^{*}_{g_{2},2}$ columns, we have $1+k_{1}^{2}=k_{2}^{2}+k_{3}^{2}$. For the $\mathcal{B}^{*}_{g_{1}}$ rows and the $\mathcal{B}^{*}_{g_{2},2}$ columns, we have
$\bm{\lambda}^{*}_{1}(\bm{\lambda}^{*}_{2,2})^{\top}+\rho^{*}_{1,2}\bm{\lambda}^{*}_{g_{1}}(\bm{\lambda}^{*}_{g_{2},2})^{\top}$ (C.5)
$=$ $\bm{\lambda}_{1}(\bm{\lambda}_{2,2})^{\top}+\rho_{1,2}\bm{\lambda}_{g_{1}}(\bm{\lambda}_{2,g^{\prime}_{2}})^{\top}$
$=$ $k_{2}(\cos{\alpha}\bm{\lambda}^{*}_{1}-\sin{\alpha}\bm{\lambda}^{*}_{g_{1}})(\bm{\lambda}^{*}_{g_{2},2})^{\top}+k_{3}(\sin{\alpha}\bm{\lambda}^{*}_{1}+\cos{\alpha}\bm{\lambda}^{*}_{g_{1}})(\bm{\lambda}^{*}_{g_{2},2})^{\top}$
$=$ $(k_{2}\cos{\alpha}+k_{3}\sin{\alpha})\bm{\lambda}^{*}_{1}(\bm{\lambda}^{*}_{g_{2},2})^{\top}+(k_{3}\cos{\alpha}-k_{2}\sin{\alpha})\bm{\lambda}^{*}_{g_{1}}(\bm{\lambda}^{*}_{g_{2},2})^{\top}.$
Then, we have $k_{1}=k_{2}\cos{\alpha}+k_{3}\sin{\alpha}$ and $\rho^{*}_{1,2}=k_{3}\cos{\alpha}-k_{2}\sin{\alpha}$, which leads to $k_{1}^{2}+(\rho^{*}_{1,2})^{2}=k_{2}^{2}+k_{3}^{2}$. Combined with $1+k_{1}^{2}=k_{2}^{2}+k_{3}^{2}$, we have $|\rho^{*}_{1,2}|=1$, which contradicts to the fact that $\Phi^{*}$ is positive definite. Thus, only the first case is allowed. Without loss of generation, we assume $g^{\prime}_{2}=g_{2}$.
For the first case, there exists some $\alpha,\beta$ such that $\bm{\lambda}_{1}=\cos{\alpha}\bm{\lambda}^{*}_{1}-\sin{\alpha}\bm{\lambda}^{*}_{g_{1}}$, $\bm{\lambda}_{g_{1}}=\sin{\alpha}\bm{\lambda}^{*}_{1}+\cos{\alpha}\bm{\lambda}^{*}_{g_{1}}$, $\bm{\lambda}_{2}=\cos{\beta}\bm{\lambda}^{*}_{2}-\sin{\beta}\bm{\lambda}^{*}_{g_{2}}$ and $\bm{\lambda}_{g_{2}}=\sin{\beta}\bm{\lambda}^{*}_{2}+\cos{\beta}\bm{\lambda}^{*}_{g_{2}}$. We then have the following equation
$\begin{pmatrix}\cos{\alpha}&\sin{\alpha}\\ -\sin{\alpha}&\cos{\alpha}\end{pmatrix}\begin{pmatrix}1&0\\ 0&\rho_{12}\end{pmatrix}\begin{pmatrix}\cos{\beta}&-\sin{\beta}\\ \sin{\beta}&\cos{\beta}\end{pmatrix}=\begin{pmatrix}1&0\\ 0&\rho^{*}_{12}\end{pmatrix},$ (C.6)
which leads to $1=\cos{\alpha}\cos{\beta}+\rho^{*}_{12}\sin{\alpha}\sin{\beta}$. Since $|\rho^{*}_{12}|<1$, we have $\cos{\alpha}\cos{\beta}=1$ and $\sin{\alpha}\sin{\beta}=0$. Without loss of generation, we assume $\cos{\alpha}=\cos{\beta}=1$. Then we have $\bm{\lambda}_{1}=\bm{\lambda}^{*}_{1}$, $\bm{\lambda}_{g_{1}}=\bm{\lambda}^{*}_{g_{1}}$, $\bm{\lambda}_{2}=\bm{\lambda}^{*}_{2}$, $\bm{\lambda}_{g_{2}}=\bm{\lambda}^{*}_{g_{2}}$ and $\rho_{12}=\rho_{12}^{*}$.
For any group $g_{3}\neq g_{1},g_{2}$, we consider the $\mathcal{B}^{*}_{g1}\cup\mathcal{B}^{*}_{g3}$ rows and $\mathcal{B}^{*}_{g1}\cup\mathcal{B}^{*}_{g3}$ columns of $\Lambda\Phi(\Lambda)^{\top}$ and $\Lambda^{*}\Phi^{*}(\Lambda^{*})^{\top}$. Similar to the the proof of $g_{2}$, there exists only one possible bi-factor structure : For some $g^{\prime}_{3}\in\{1,\ldots,G\}$, we have
$\begin{pmatrix}\bm{\lambda}_{1}&\bm{\lambda}_{g_{1}}&\mathbf{0}\\ \bm{\lambda}_{3}&\mathbf{0}&\bm{\lambda}_{g^{\prime}_{3}}\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&\rho_{1,3}\\ 0&\rho_{1,3}&1\end{pmatrix}\begin{pmatrix}\bm{\lambda}_{1}^{\top}&\bm{\lambda}_{3}^{\top}\\ \bm{\lambda}_{g_{1}}^{\top}&\mathbf{0}^{\top}\\ \mathbf{0}^{\top}&\bm{\lambda}_{g^{\prime}_{3}}^{\top}\end{pmatrix}=\begin{pmatrix}\bm{\lambda}^{*}_{1}&\bm{\lambda}^{*}_{g_{1}}&\mathbf{0}\\ \bm{\lambda}^{*}_{3}&\mathbf{0}&\bm{\lambda}^{*}_{g_{3}}\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&\rho_{1,3}^{*}\\ 0&\rho_{1,3}^{*}&1\end{pmatrix}\begin{pmatrix}(\bm{\lambda}_{1}^{*})^{\top}&(\bm{\lambda}_{3}^{*})^{\top}\\ (\bm{\lambda}_{g_{1}}^{*})^{\top}&\mathbf{0}^{\top}\\ \mathbf{0}^{\top}&(\bm{\lambda}_{g_{3}}^{*})^{\top},\end{pmatrix}$ (C.7)
where $\bm{\lambda}_{i}=\Lambda[\mathcal{B}^{*}_{g_{i}},\{1\}]$, $\bm{\lambda}_{g^{\prime}_{i}}=\Lambda[\mathcal{B}^{*}_{g_{i}},\{1+g^{\prime}_{i}\}]$, $\bm{\lambda}^{*}_{i}=\Lambda^{*}[\mathcal{B}^{*}_{g_{i}},\{1\}]$ and $\bm{\lambda}^{*}_{g_{i}}=\Lambda^{*}[\mathcal{B}^{*}_{g_{i}},\{1+g_{1}\}]$ for $i=1,3$. $\rho_{1,3}=\phi_{1+g^{\prime}_{1},1+g^{\prime}_{3}}$ and $\rho^{*}_{1,3}=\phi^{*}_{1+g_{1},1+g_{3}}$. We then have
$\bm{\lambda}^{*}_{1}(\bm{\lambda}^{*}_{3})^{\top}+\rho^{*}_{1,3}\bm{\lambda}^{*}_{g_{1}}(\bm{\lambda}^{*}_{g_{3}})^{\top}=\bm{\lambda}_{1}(\bm{\lambda}_{3})^{\top}+\rho_{1,3}\bm{\lambda}_{g_{1}}(\bm{\lambda}_{g^{\prime}_{3}})^{\top}$ (C.8)
$\bm{\lambda}^{*}_{3}(\bm{\lambda}^{*}_{3})^{\top}+\bm{\lambda}^{*}_{g_{3}}(\bm{\lambda}^{*}_{g_{3}})^{\top}=\bm{\lambda}_{3}(\bm{\lambda}_{3})^{\top}+\bm{\lambda}_{g_{3}}(\bm{\lambda}_{g_{3}})^{\top}.$
Since we have proved $\bm{\lambda}^{*}_{1}=\bm{\lambda}_{1}$ and $\bm{\lambda}^{*}_{g_{1}}=\bm{\lambda}_{g_{1}}$, we then have $\bm{\lambda}^{*}_{3}=\bm{\lambda}_{3}$, $\bm{\lambda}^{*}_{g_{3}}=\bm{\lambda}_{g^{\prime}_{3}}$ or $\bm{\lambda}^{*}_{g_{3}}=-\bm{\lambda}_{g^{\prime}_{3}}$ and $\rho^{*}_{1,3}=\rho_{1,3}$ or $\rho^{*}_{1,3}=-\rho_{1,3}$. Then, the proof is completed.
C.2 Identifiability of Estimated Bi-factor Structure in Real Data Example
For any extended bi-factor model, we denote by $\mathcal{Q}_{g}=\{i\in\mathcal{B}_{g}|\lambda_{i,g+1}\neq 0\}$ and $H=\{g|\Lambda[\mathcal{B}_{g},\{1,1+g\}]\mbox{~{}has column rank 2}\}$. The following condition is the necessary condition for the identifiability of the extended bi-factor model under a known bi-factor structure, as proposed in Theorem 3 of Fang et al., (2021).
Condition 4.
$|H|\geq 2$; $|\mathcal{Q}_{g}|\geq 2$ for $1,\ldots,G$ and $|\mathcal{Q}_{g}|\geq 3$ for $g$ that satisfies $\phi_{1+g,1+g^{\prime}}=0$ for all $g^{\prime}\neq g$.
We then propose the following condition for the identifiability of parameters when the true bi-factor structure is the same as the estimated structure in Section 4.
Condition 5.
$|H^{*}|\geq 3$. In addition, there exist $g1,g2,g3\in H^{*}$ such that (1) $|Q_{g1}^{*}|\geq 3$, $|Q_{g2}^{*}|\geq 3$ and $|Q_{g3}^{*}|\geq 3$, (2) any 2 rows of $\Lambda^{*}[\mathcal{B}_{g1}^{*},\{1+g_{1}\}]$ are linearly independent.
Remark 4.
Condition 5 is easy to check in practice and $\widehat{\Lambda}$ in Section 4 satisfies Condition 5.
Theorem 2.
Suppose the true bi-factor structure follows $\widehat{\Lambda}$ in Section 4. Let $\Lambda^{*}$, $\Phi^{*}$ and $\Psi^{*}$ be the true parameters such that Condition 4and 5 are satisfied. For any parameters $\Lambda$, $\Phi$ and $\Psi$ that satisfy Condition 4 and $\Lambda^{*}\Phi^{*}(\Lambda^{*})^{\top}+\Psi^{*}=\Lambda\Phi(\Lambda)^{\top}+\Psi$, there exists a diagonal sign-flip matrix $D\in\mathcal{D}$ and a permutation matrix $P\in\mathcal{P}$ such that $\Lambda=\Lambda^{*}PD$, $\Phi=DP^{\top}\Phi^{*}PD$ and $\Psi^{*}=\Psi$.
Proof of Theorem 2 : Without loss of generation, we assume groups $1,2,3$ of $\Lambda^{*}$ satisfy Condition 5 and $|\mathcal{B}_{5}^{*}|=|\mathcal{B}_{6}^{*}|=|\mathcal{B}_{7}^{*}|=2$. The proof consists of three part: (1) show $\psi_{i}=\psi^{*}_{i}$ for $i\in\mathcal{B}_{1},\ldots,\mathcal{B}_{4}$;(2) show there exists some $D\in\mathcal{D}$ and $P\in\mathcal{P}$ such that $\Lambda[\cup_{i=1}^{4}\mathcal{B}^{*}_{i},:]=\Lambda^{*}[\cup_{i=1}^{4}\mathcal{B}^{*}_{i},:]PD$ and $(PD\Phi DP^{\top})[\{1,\ldots,5\},\{1,\ldots,5\}]=\Phi^{*}[\{1,\ldots,5\},\{1,\ldots,5\}]$. (3) show $\Lambda[\cup_{i=5}^{7}\mathcal{B}^{*}_{i},:]=\Lambda^{*}[\cup_{i=5}^{7}\mathcal{B}^{*}_{i},:]PD$, $(PD\Phi DP^{\top})[\{6,7,8\},\{6,7,8\}]=\Phi^{*}[\{6,7,8\},\{6,7,8\}]$ and $\psi_{i}=\psi_{i}^{*}$ for $i\in\mathcal{B}_{5},\ldots,\mathcal{B}_{7}$.
We first show that $\psi_{i}=\psi^{*}_{i}$ for $i\in\mathcal{B}_{1},\ldots,\mathcal{B}_{4}$. According to the proof of Theorem 5.1 of Anderson and Rubin, (1956), for any $j\in\{1,\ldots,J\}$, if the $j$th row $\Lambda^{*}$ is deleted and there exist two disjoint submatrices of the rest of $\Lambda^{*}$ with full columns rank, then $\psi_{j}=\psi_{j}^{*}$. We now construct the partitions for $j\in\cup_{i=1}^{4}\mathcal{B}^{*}_{i}$. By Condition 5, there exist 3 different observations $i_{k1},i_{k2},i_{k3}\in\mathcal{B}_{k}^{*}$ such that $\Lambda^{*}[\{i_{k1}\},\{1+k\}]$,$\Lambda^{*}[\{i_{k2}\},\{1+k\}]$ and $\Lambda^{*}[\{i_{k3}\},\{1+k\}]\neq 0$ for $k=1,2,3,4$. Furthermore, there exist 2 different observations $i_{k4},i_{k5}\in\mathcal{B}_{k}^{*}$ such that $\Lambda^{*}[\{i_{k4},i_{k5}\},\{1,1+k\}]$ has rank 2 for $k=1,2,3$. By Condition 4, there exist 2 different observations $\Lambda^{*}[\{i_{k2}\},\{1+k\}]$ and $\Lambda^{*}[\{i_{k2}\},\{1+k\}]\neq 0$ for $k=5,6,7$.
Now for $j\in\mathcal{B}^{*}_{4}$, for example $j=i_{4,3}$, we select the first submatrix that consists of rows $i_{14},i_{15},i_{21},i_{31},\ldots,\\ i_{71}$ of $\Lambda^{*}$ and the second submatrix that consists of rows $i_{11},i_{24},i_{25},i_{32},\ldots,i_{72}$, where $i_{11}\neq i_{14},i_{15}$ and $i_{21}\neq i_{24},i_{25}$. It is easy to check that the two submatrices have column rank 8. For $j\in\mathcal{B}_{1}^{*}$, for example $j=i_{11}$, we construct the first submatrix that consists of rows $i_{12},i_{24},i_{25},i_{31},i_{41},\ldots,i_{71}$ of $\Lambda^{*}$, where $i_{31}\neq i_{34},i_{35}$ and the second submatrix that consists of rows $i_{13},i_{21},i_{34},i_{35},i_{42},\ldots,i_{72}$ of $\Lambda^{*}$, where $i_{21}\neq i_{24},i_{25}$. It is easy to check that the two submatrices have column rank 8. The construction of submatrices for $j\in\mathcal{B}^{*}_{2},\mathcal{B}^{*}_{3}$ is similar to the case of $j\in\mathcal{B}^{*}_{1}$. Then, by Theorem 5.1 of Anderson and Rubin, (1956), $\psi_{i}=\psi^{*}_{i}$ for $i\in\mathcal{B}_{1},\ldots,\mathcal{B}_{4}$.
Secondly, with a similar argument in Theorem 1, there exists some $D\in\mathcal{D}$ and $P\in\mathcal{P}$ such that $\Lambda[\cup_{i=1}^{4}\mathcal{B}^{*}_{i},:]=\Lambda^{*}[\cup_{i=1}^{4}\mathcal{B}^{*}_{i},:]PD$ and $(PD\Phi DP^{\top})[\{1,\ldots,5\},\{1,\ldots,5\}]=\Phi^{*}[\{1,\ldots,5\},\{1,\ldots,5\}]$. Without loss of generation, we assume $P$ and $D$ are identity matrices.
Thirdly, we denote by $\mathcal{B}_{5}^{*}=\{i_{5},i_{6}\}$, $\mathcal{B}_{6}^{*}=\{i_{7},i_{8}\}$ and $\mathcal{B}_{7}^{*}=\{i_{9},i_{10}\}$. Since $\Lambda$, $\Phi$ and $\Psi$ satisfy Condition 4, there exists three types of possible of bi-factor structure of $\mathcal{B}_{5}$, $\mathcal{B}_{6}$ and $\mathcal{B}_{7}$: (1)$\mathcal{B}_{5}=\{i_{5},i_{6}\}$, $\mathcal{B}_{6}=\{i_{7},i_{8}\}$ and $\mathcal{B}_{7}=\{i_{9},i_{10}\}$; (2) $\mathcal{B}_{5}=\{i_{5},i_{10}\}$, $\mathcal{B}_{6}=\{i_{6},i_{7}\}$ and $\mathcal{B}_{7}=\{i_{8},i_{9}\}$; (3) $\mathcal{B}_{5}=\{i_{5},i_{6}\}$, $\mathcal{B}_{6}=\{i_{7},i_{9}\}$ and $\mathcal{B}_{7}=\{i_{8},i_{10}\}$.
We now show that the second case is not allowed. We consider the off diagonal part of $\mathcal{B}_{1}^{*}$, $\mathcal{B}_{5}^{*}$, $\mathcal{B}_{6}^{*}$ and $\mathcal{B}_{7}^{*}$ rows and columns of $\Lambda^{*}\Phi^{*}(\Lambda^{*})^{\top}+\Psi^{*}$ and $\Lambda\Phi(\Lambda)^{\top}+\Psi$. Firstly, we have $\Lambda^{*}[\mathcal{B}^{*}_{1},\{1\}]\lambda^{*}_{i_{5},1}+\phi^{*}_{2,6}\Lambda^{*}[\mathcal{B}^{*}_{1},\{2\}]\lambda^{*}_{i_{5},6}=\Lambda[\mathcal{B}^{*}_{1},\{1\}]\lambda_{i_{5},1}+\phi_{2,6}\Lambda[\mathcal{B}^{*}_{1},\{2\}]\lambda_{i_{5},6}$. Since $\Lambda[\mathcal{B}^{*}_{1},\{1,2\}]$ are linear independent and we have proved $\Lambda[\mathcal{B}^{*}_{1},\{1\}]=\Lambda[\mathcal{B}_{1},\{1\}]$ and $\Lambda[\mathcal{B}^{*}_{1},\{2\}]=\Lambda[\mathcal{B}_{1},\{2\}]$, we have $\lambda^{*}_{i_{5},1}=\lambda_{i_{5},1}$. Similarly, we have $\lambda^{*}_{i_{k},1}=\lambda_{i_{k},1}$ for $k=5,\ldots,10$. Next, we have $\lambda^{*}_{i_{5},1}\lambda^{*}_{i_{6},1}+\lambda^{*}_{i_{5},6}\lambda^{*}_{i_{6},6}=\lambda_{i_{5},1}\lambda_{i_{6},1}+\phi_{6,7}\lambda_{i_{5},6}\lambda_{i_{6},7}$, which leads to $\phi_{6,7}=\lambda^{*}_{i_{5},6}\lambda^{*}_{i_{6},6}/(\lambda_{i_{5},6}\lambda_{i_{6},7})$. We can similarly have $\phi_{67}=\lambda^{*}_{i_{5},6}\lambda^{*}_{i_{6},6}/(\lambda_{i_{5},6}\lambda_{i_{6},7})=\phi^{*}_{6,7}\lambda^{*}_{i_{5},6}\lambda^{*}_{i_{7},7}/(\lambda_{i_{5},6}\lambda_{i_{7},7})=\phi^{*}_{6,8}\lambda^{*}_{i_{6},6}\lambda^{*}_{i_{10},8}/(\lambda_{i_{6},7}\lambda_{i_{10},6})=\phi^{*}_{7,8}\lambda^{*}_{i_{7},7}\lambda^{*}_{i_{10},8}/(\lambda_{i_{7},7}\lambda_{i_{10},6})$. Combing these equations, we have $\phi^{*}_{6,7}\phi^{*}_{6,8}=\phi^{*}_{7,8}$. Also, by Condition 4, these equations lead to $\phi^{*}_{6,7},\phi^{*}_{6,8},\phi^{*}_{7,8}\neq 0$. In the same manner, we have $\phi^{*}_{6,8}\phi^{*}_{7,8}=\phi^{*}_{6,7}$ and $\phi^{*}_{6,7}\phi^{*}_{7,8}=\phi^{*}_{6,8}$. Then, we have $|\phi^{*}_{6,7}|=|\phi^{*}_{6,8}|=|\phi^{*}_{7,8}|=1$, which violates the assumption that $\Phi^{*}$ is positive definite.
Next, we show that the third case is not allowed. Similar to the proof in the second case, we first have $\lambda^{*}_{i_{k},1}=\lambda_{i_{k},1}$ for $k=5,\ldots,10$. Then we have the following 6 equations:
$\lambda^{*}_{i_{7},7}\lambda^{*}_{i_{8},7}=\phi_{7,8}\lambda_{i_{7},7}\lambda_{i_{8},8},$ (C.9)
$\phi^{*}_{7,8}\lambda^{*}_{i_{7},7}\lambda^{*}_{i_{9},8}=\lambda_{i_{7},7}\lambda_{i_{9},7},$
$\phi^{*}_{7,8}\lambda^{*}_{i_{7},7}\lambda^{*}_{i_{10},8}=\phi_{7,8}\lambda_{i_{7},7}\lambda_{i_{10},8},$
$\phi^{*}_{7,8}\lambda^{*}_{i_{8},7}\lambda^{*}_{i_{9},8}=\phi_{7,8}\lambda_{i_{8},8}\lambda_{i_{9},7},$
$\phi^{*}_{7,8}\lambda^{*}_{i_{8},7}\lambda^{*}_{i_{10},8}=\lambda_{i_{8},8}\lambda_{i_{10},8},$
$\lambda^{*}_{i_{9},8}\lambda^{*}_{i_{10},8}=\phi_{7,8}\lambda_{i_{9},7}\lambda_{i_{10},8}.$
With Condition 4, according to the first equation above, we have $\phi_{7,8}\neq 0$. By the 6 equations, we also have $(\phi^{*}_{7,8})^{4}\phi_{7,8}^{4}=\phi_{7,8}^{2}$. Then we have $|\phi^{*}_{7,8}|=|\phi_{7,8}|=1$, which violates the assumption that $\Phi^{*}$ is positive definite.
Thus, only the case such that $\mathcal{B}_{5}=\{i_{5},i_{6}\}$, $\mathcal{B}_{6}=\{i_{7},i_{8}\}$ and $\mathcal{B}_{7}=\{i_{9},i_{10}\}$ is possible. Similar to the proof before, we have $\lambda^{*}_{i_{k},1}=\lambda_{i_{k},1}$ for $k=5,\ldots,10$. Similar to the proof of Proposition 1 of Fang et al., (2021), by Condition 4, there exists some $g\in\{1,\ldots,7\}$ such that $\phi_{6,1+g}\neq 0$, we then have $c(\lambda^{*}_{i_{5},6},\lambda^{*}_{i_{6},6})=(\lambda_{i_{5},6},\lambda_{i_{6},6})$ for some constant $c$. Since $\lambda^{*}_{i_{5},6}\lambda^{*}_{i_{6},6}=\lambda_{i_{5},6}\lambda_{i_{6},6}$, we have $c=1$ or $-1$, which leads to $(\lambda^{*}_{i_{5},6},\lambda^{*}_{i_{6},6})=(\lambda_{i_{5},6},\lambda_{i_{6},6})$ or $(\lambda^{*}_{i_{5},6},\lambda^{*}_{i_{6},6})=-(\lambda_{i_{5},6},\lambda_{i_{6},6})$. Then, we further have $\Psi_{i_{5}}=\Psi^{*}_{i_{5}}$ and $\Psi_{i_{6}}=\Psi^{*}_{i_{6}}$. The proof for observations $i_{7},\ldots,i_{10}$ is similar. Then, the proof is completed.